OpenStudy (anonymous):

Joey and Mark go on vacation and travel in opposite directions. Joey leaves the house at 8:00AM and averages 35mph. Mark leaves at 9:00AM and averages 40mph. At what time will they be 225 miles apart?

OpenStudy (kropot72):

Distance traveled by Joey in the one hour before Mark starts is 35 miles. Therefore the required extra separation distance from when Mark starts is 225 - 35 = 190 miles. The separation speed when both Joey and Mark are traveling is 35 + 40 = 75 mph. The time needed to travel 190 miles at 75 mph is 190/75 hours. Therefore the time at which they are 225 miles apart is (9 + 190/75)AM. Now you need to convert 190/75 hours into hours and minutes. Can you do that?

OpenStudy (anonymous):

No,I'm not sure :/

OpenStudy (kropot72):

\[\large \frac{190}{75}=2\frac{40}{75}=2\frac{8}{15}\ hours\] \[\large \frac{8}{15}\times\frac{60}{1}=32\ minutes\] Now to find the time at which they are 225 miles apart, you need to add 2 hours 32 minutes to 9.00AM.

OpenStudy (kropot72):

@Zoexo Have you got it yet?